OpenStudy (anonymous):

f(x)={x^2+2x x<=-1} {x+8 x>-1} find f(-4) f(-1)

jimthompson5910 (jim_thompson5910):

let's find f(-4) first

jimthompson5910 (jim_thompson5910):

f(-4) means find f(x) when x = -4

OpenStudy (anonymous):

I solved f(6) and got 14.. but got the others wrong.

jimthompson5910 (jim_thompson5910):

what did you get for f(6)?

OpenStudy (anonymous):

my answer is 14 and I got it right.

jimthompson5910 (jim_thompson5910):

keep in mind that when x = -4, we're looking at the first piece (since x <= -1 here)

jimthompson5910 (jim_thompson5910):

f(6) = 14 is correct

OpenStudy (anonymous):

so I just plug in right?

jimthompson5910 (jim_thompson5910):

because when x = 6, which is greater than -1, we look at the corresponding piece x+8

jimthompson5910 (jim_thompson5910):

but when x = -4, we look at the first piece since f(x) = x^2 + 2x when x <= -1

OpenStudy (anonymous):

-24?

jimthompson5910 (jim_thompson5910):

f(x) = x^2 + 2x f(-4) = (-4)^2 + 2(-4) f(-4) = 16 + 2(-4) <<--- This is where you made your mistake f(-4) = 16 - 8 f(-4) = 8

jimthompson5910 (jim_thompson5910):

(-4)^2 = 16 and NOT (-4)^2 = -16

OpenStudy (anonymous):

gahh, okay.

jimthompson5910 (jim_thompson5910):

that's ok, it gets a lot of people

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome